# Two-Timescale Design for Movable Antennas Enabled-Multiuser MIMO Systems

### v1

@@ IV-A Ergodic Rate Analysis with ZF beamforming
×$\!\!\!\times\!\!\frac{(\boldsymbol{I}_{N}-\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)}{\lVert(\boldsymbol{I}_{N}-\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)\rVert},\!$ (32)
where $p_{m}$ represents the power equally allocated to user $m$, and $\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)$ is defined as the projection matrix for user $m$’s channel, ensuring orthogonality to other users’ channels
$\!\!\!\!\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)\!=\!\left[\boldsymbol{h}_{1}\left(\boldsymbol{t}\right),\dots,\boldsymbol{h}_{m-1}\left(\boldsymbol{t}\right),\boldsymbol{h}_{m+1}\left(\boldsymbol{t}\right),\dots,\boldsymbol{h}_{M}\left(\boldsymbol{t}\right)\right].\!\!\!$ (33)
@@
×$\!\times\!\frac{|\boldsymbol{h}_{m}^{H}\left(\boldsymbol{t}\right)(\boldsymbol{I}_{N}\!-\!\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\!\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)|^{2}}{\lVert(\boldsymbol{I}_{N}-\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)\rVert^{2}}$
@@
[…]
The exact expression for the ergodic rate under ZF beamforming is challenging to handle. To address this, we employ Jensen’s inequality to derive a tight lower-bound approximation [33], which is expressed as
@@ IV-B Antenna Position Optimization Based on AO Method
Despite having closed-form expressions for the ergodic rate in (37), the antenna positions remain intricately coupled in the implicit function $[\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}\right)]_{mm}$, which depends primarily on the LoS components. To tackle this complexity, we propose an element-wise antenna position optimization framework that leverages Woodbury’s identity, the MM, and the SCA method.
@@
The inequality holds strictly with MAs, whereas equality is achieved in the FPA case for the $n$-th antenna. This observation demonstrates that MAs offer a performance gain over FPAs by reducing the inverse channel matrix component $[\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}_{n}\right)]_{mm}$. Specifically, this reduction indicates a decrease in channel correlation between user $m$ and other users while maintaining the channel power gain for user $m$.
Then, $[\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}_{n}\right)]_{mm}$ can be expressed more compactly as
@@
$\boldsymbol{l}_{n,m}^{H}=\left[\boldsymbol{\varTheta}_{2,n}^{-1}\boldsymbol{\varLambda}_{2}\right]_{\left(m,:\right)}m$ (47b)
@@
with $[\cdot]_{\left(m,:\right)}$ referred to the $m$-th row vector of a matrix.
@@ IV-C Constructing a Surrogate Function via MM and SCA
$R_{\text{ZF},m}^{\text{lb},1}\left(\boldsymbol{t}_{n}\right)\leq R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)$
@@
×$\,\,\times\!\big{\{}2\lambda_{\max}\left(\boldsymbol{X}_{n,m}\right)M-\boldsymbol{\bar{g}}_{n}^{H}\left(\boldsymbol{t}_{n}^{\ell}\right)\boldsymbol{X}_{n,m}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}^{\ell}\right)$
@@
$\,\,\Big{(}\frac{\partial\big{(}\frac{\boldsymbol{\bar{g}}_{n}^{H}\left(\boldsymbol{t}_{n}\right)\boldsymbol{Y}_{n}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)}{r_{n,m}\left(\boldsymbol{t}_{n}\right)}\big{)}}{\partial\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)}\bigg{|}_{\boldsymbol{t}_{n}=\boldsymbol{t}_{n}^{\ell}}\Big{)}^{H}\left(\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)-\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}^{\ell}\right)\right).\!\!\!\!$ (52)
@@
$R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)$ $=\log\big{(}1+\eta_{m}\big{(}\chi_{n,m}^{\ell}+\text{Re}\left\{\boldsymbol{q}_{n,m}^{T}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)\right\}\big{)}\big{)}$
$=\log\Big{(}1+\eta_{m}\big{(}\chi_{n,m}^{\ell}+\sum_{u=1}^{M}{\left|q_{n,m,u}^{\ell}\right|}$
×$\quad\quad\quad\quad\quad\times\cos(\frac{2\pi}{\lambda}\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{u}-\angle q_{n,m,u}^{\ell})\big{)}\Big{)},$ (55)
@@
$R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)$ $\leq R_{\text{ZF},m}^{\text{lb},3}\left(\boldsymbol{t}_{n}\right)$
$=\log_{2}\Big{(}1+\eta_{m}\big{(}\chi_{n,m}^{\ell}+F_{n,m}(\boldsymbol{t}_{n}^{\ell})+\nabla z_{m}(\boldsymbol{t}_{n}^{\ell})^{T}$
×$\quad\quad\quad\times(\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell})-\frac{\psi_{m,n}}{2}\|\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\|^{2}\big{)}\Big{)},$ (59)
@@ IV-D Overall Algorithm and Computational Complexity Analysis
4:     For each iteration, solve the problem (P3.n) for $N$ antennas using convex optimization methods obtain updated antenna positions $\boldsymbol{t}_{n}^{\ell+1}$.
@@
The overall AO-based algorithm is detailed in Algorithm 2, which iteratively solves (P5.n) to provide a solution for the problem (P4). The algorithm guarantees convergence to a locally optimal solution, as the objective function value is non-decreasing in each iteration. Since the problem is a convex QCP, the overall computational complexity of Algorithm 2 is $\mathcal{O}(L_{2}\ln\frac{1}{\varepsilon}N^{4.5})$, where $L_{2}$ represents the number of iterations until convergence, and $\varepsilon$ denotes the prescribed accuracy.
V Simulation Results

### v2

@@ IV-A Ergodic Rate Analysis with ZF beamforming
$\!\!\!\times\!\!\frac{(\boldsymbol{I}_{N}-\boldsymbol{B}_{m}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{m}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)}{\lVert(\boldsymbol{I}_{N}-\boldsymbol{B}_{m}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{m}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)\rVert},\!$ (32)
where $p_{m}$ represents the power equally allocated to user $m$, and $\boldsymbol{B}_{m}\left(\boldsymbol{t}\right)$ is defined as the projection matrix for user $m$’s channel, ensuring orthogonality to other users’ channels
$\!\!\!\!\boldsymbol{B}_{m}\left(\boldsymbol{t}\right)\!=\!\left[\boldsymbol{h}_{1}\left(\boldsymbol{t}\right),\dots,\boldsymbol{h}_{m-1}\left(\boldsymbol{t}\right),\boldsymbol{h}_{m+1}\left(\boldsymbol{t}\right),\dots,\boldsymbol{h}_{M}\left(\boldsymbol{t}\right)\right].\!\!\!$ (33)
@@
$\!\times\!\frac{|\boldsymbol{h}_{m}^{H}\left(\boldsymbol{t}\right)(\boldsymbol{I}_{N}\!-\!\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\!\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)|^{2}}{\lVert(\boldsymbol{I}_{N}-\boldsymbol{B}_{n}\left(\boldsymbol{t}\right)(\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{B}_{n}\left(\boldsymbol{t}\right))^{-1}\boldsymbol{B}_{m}^{H}\left(\boldsymbol{t}\right))\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)\rVert^{2}}$
@@
$\text{s}.\text{t}.$ $\,\,(6\text{c}),(6\text{d}).$
The exact expression for the ergodic rate under ZF beamforming is challenging to handle. To address this, we employ Jensen’s inequality to derive a tight lower-bound approximation [33], which is expressed as [^6] Due to the nature of ZF beamforming that completely cancels multiuser interference, the proposed design cannot be directly applicable when $M>N$. However, we can adopt user grouping to divide and select smaller subsets of users. In this way, MA still offers extra degrees of freedom by shaping channel conditions and enhancing ergodic rate.
@@
Remark 3: For the ZF beamforming that eliminates multiuser interference, the inverse channel term $(\boldsymbol{H}^{H}\boldsymbol{H})^{-1}]_{mm}$ determines the SNR and thus the ergodic rate. From a statistical perspective, the sum-rate performance depends strongly on the ratio $\frac{N}{M}$ and the SNR [41]. As $M$ increases and approaches $N$, spatial degrees of freedom diminish, making the channel matrix ill-conditioned [42]. This amplifies $(\boldsymbol{H}^{H}\boldsymbol{H})^{-1}]_{mm}$ and reduces the effective SNR for each user. Hence. FPA-ZF systems experience a marked decline in sum rate due to limited channel diversity and higher susceptibility to correlation. In the MA-enabled ZF system, antenna repositioning offers an additional degree of freedom to reshape $\boldsymbol{H}(\boldsymbol{t})$. Specifically, in the two-timescale framework, optimizing $\boldsymbol{t}$ to minimize $\left[\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}\right)\right]_{mm}$ can reduce off-diagonal elements in $\boldsymbol{H}^{H}\left(\boldsymbol{t}\right)\boldsymbol{H}\left(\boldsymbol{t}\right)$ (i.e., channel correlations), thus improving channel orthogonality. At the same time, the diagonal elements (overall channel power) remain largely preserved and thus increase the ergodic rate.
@@ IV-B Antenna Position Optimization Based on AO Method
Despite having closed-form expressions for the ergodic rate in (37), the antenna positions remain intricately coupled in the implicit function $\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}\right)]_{mm}$, which depends primarily on the LoS components. To tackle this complexity, we propose an element-wise antenna position optimization framework that leverages Woodbury’s identity, the MM, and the SCA method.
@@
The inequality holds strictly with MAs, whereas equality is achieved in the FPA case for the $n$-th antenna. This observation demonstrates that MAs offer a performance gain over FPAs by reducing the inverse channel matrix component $\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}_{n}\right)]_{mm}$. Specifically, this reduction indicates a decrease in channel correlation between user $m$ and other users while maintaining the channel power gain for user $m$.
Then, $\boldsymbol{\varSigma}^{-1}\left(\boldsymbol{t}_{n}\right)]_{mm}$ can be expressed more compactly as
@@
$\boldsymbol{l}_{n,m}^{H}=\left[\boldsymbol{\varTheta}_{2,n}^{-1}\boldsymbol{\varLambda}_{2}\right]_{\left(m,:\right)},$ (47b)
@@
with $\cdot]_{\left(m,:\right)}$ referred to the $m$-th row vector of a matrix.
@@ IV-C Constructing a Surrogate Function via MM and SCA
$R_{\text{ZF},m}^{\text{lb},1}\left(\boldsymbol{t}_{n}\right)\geq R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)$
@@
$\,\,\times\!\big\{2\lambda_{\max}\left(\boldsymbol{X}_{n,m}\right)M-\boldsymbol{\bar{g}}_{n}^{H}\left(\boldsymbol{t}_{n}^{\ell}\right)\boldsymbol{X}_{n,m}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}^{\ell}\right)$
@@
$+\Big(\frac{\partial\big(\frac{\boldsymbol{\bar{g}}_{n}^{H}\left(\boldsymbol{t}_{n}\right)\boldsymbol{Y}_{n}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)}{r_{n,m}\left(\boldsymbol{t}_{n}\right)}\big)}{\partial\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)}\bigg|_{\boldsymbol{t}_{n}=\boldsymbol{t}_{n}^{\ell}}\Big)^{H}\left(\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)-\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}^{\ell}\right)\right).\!\!\!\!$ (52)
@@
$R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)$ $=\log_{2}\big(1+\eta_{m}\big(\chi_{n,m}^{\ell}+\text{Re}\left\{\boldsymbol{q}_{n,m}^{T}\boldsymbol{\bar{g}}_{n}\left(\boldsymbol{t}_{n}\right)\right\}\big)\big)$
$=\log_{2}\Big(1+\eta_{m}\big(\chi_{n,m}^{\ell}+\sum_{u=1}^{M}{\left|q_{n,m,u}^{\ell}\right|}$
$\quad\quad\quad\quad\quad\times\cos(\frac{2\pi}{\lambda}\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{u}-\angle q_{n,m,u}^{\ell})\big)\Big),$ (55)
@@
$R_{\text{ZF},m}^{\text{lb},2}\left(\boldsymbol{t}_{n}\right)\geq R_{\text{ZF},m}^{\text{lb},3}\left(\boldsymbol{t}_{n}\right)=\log_{2}\big(1+\eta_{m}\big(\chi_{n,m}^{\ell}+F_{n,m}(\boldsymbol{t}_{n}^{\ell})$
$+\nabla F_{m}(\boldsymbol{t}_{n}^{\ell})^{T}\times(\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell})-\frac{\psi_{m,n}}{2}\|\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\|^{2}\big)\big),$ (59)
@@ IV-D Overall Algorithm and Computational Complexity Analysis
4:  For each iteration, solve the problem (P3.n) for NN antennas using convex optimization methods obtain updated antenna positions $\boldsymbol{t}_{n}^{\ell+1}$.
@@
The overall AO-based algorithm is detailed in Algorithm 2, which iteratively solves (P5.n) to provide a solution to the problem (P4). The algorithm guarantees convergence to a locally optimal solution, as the objective function value is non-decreasing in each iteration. Algorithm 2 optimizes the MA positions under ZF beamforming, extending the AO and SCA framework of MA-MRT with an additional MM step, which significantly increases complexity. The total computational complexity rises to $\mathcal{O}(L_{2}(N^{2.5}\log\frac{1}{\varepsilon}+NM^{4}))$, mainly originating from tcomputing the surrogate function that involves matrix inverses and eigenvalue calculations in $\mathcal{O}(M^{3})$ for each user and each antenna in an AO iteration, totaling $\mathcal{O}(NM^{4})$. Apart from this, the QCP solving step is identical to Algorithm 1. In the small timescale, beamforming vectors are computed using closed-form solutions once antenna positions are fixed. MRT beamforming requires $\mathcal{O}(MN)$ operations to compute the beamforming vectors based on channel coefficients, while the ZF beamforming involves a matrix inversion, costing $\mathcal{O}(M^{3})$. These low-complexity operations are executed per channel realization, enabling efficient adaptation to real-time channel variations.
Although large-timescale optimization has a complexity of $\mathcal{O}(L_{1}(N^{2.5}\log\frac{1}{\varepsilon}+NM^{2}+MN^{2}))$ for MA-MRT and $\mathcal{O}(L_{2}(N^{2.5}\log\frac{1}{\varepsilon}+NM^{4}))$ for MA-ZF, for MA-ZF, its computational burden is alleviated by the temporal separation inherent in the two-timescale framework. This framework improves practical feasibility by distributing computational tasks across different timescales, balancing performance optimization with efficiency. Specifically, antenna position optimization operates on a large timescale (e.g., every 10 seconds), covering multiple time slots where user distributions and channel statistics remain relatively stable. Meanwhile, small-timescale operations (e.g., every 10 milliseconds) manage real-time channel variations through conventional channel estimation and beamforming. By significantly reducing the frequency of computationally intensive antenna position updates, this separation makes the proposed MA-enabled system both efficient and scalable.
V Ergodic Rate with Spatial Correlation
In this section, we justify our modeling assumptions for the NLoS channel components by extending the analysis beyond the simplified i.i.d. framework. We derive more accurate closed-form expressions for the ergodic rate in the MA-enabled MU-MIMO system under two-timescale design, incorporating a spatial correlation matrix $\boldsymbol{S}(\boldsymbol{t})$ that captures the covariance structure of the NLoS components. These derivations are provided for both MRT and ZF beamforming.
With the spatial correlation matrix $\boldsymbol{S}(\boldsymbol{t})$, the Rician channel model is restated as follows
$\boldsymbol{h}_{m}\left(\boldsymbol{t}\right)=\sqrt{\frac{\kappa_{m}\beta_{m}}{\kappa_{m}+1}}\boldsymbol{\bar{h}}_{m}\left(\boldsymbol{t}\right)+\sqrt{\frac{\beta_{m}}{\kappa_{m}+1}}\boldsymbol{S}^{\frac{1}{2}}\left(\boldsymbol{t}\right)\boldsymbol{u}_{m},$ (69)
[…]
$\!\!R_{\text{MRT},m}^{app}\!=\!\log_{2}\Bigg(\!1\!+\!\frac{\beta_{m}^{2}\Big(\frac{2\kappa_{m}\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{S}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}\left(\boldsymbol{t}\right)+\text{tr}\left(\boldsymbol{S}^{2}\left(\boldsymbol{t}\right)\right)}{\left(\kappa_{m}+1\right)^{2}}+N^{2}\Big)}{\sum_{j=1,j\neq m}^{M}{\beta_{m}\beta_{j}\frac{\kappa_{m}\kappa_{j}\left|\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\right|^{2}+\kappa_{m}\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\boldsymbol{S}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}\left(\boldsymbol{t}\right)+\kappa_{j}\boldsymbol{\bar{h}}_{j}^{H}\left(\boldsymbol{t}\right)\boldsymbol{S}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)\left.+\text{tr}\left(\boldsymbol{S}^{2}\left(\boldsymbol{t}\right)\right)\right)}{\left(\kappa_{m}+1\right)\left(\kappa_{j}+1\right)}\!+\!\frac{\sigma_{m}^{2}}{P}N\!\sum_{m=1}^{M}{\beta_{m}}}}\!\Bigg)\!.\!\!\!$ (70)
$\!\!\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
For ZF beamforming, following (35) and (37), to further derive the closed-form ergodic rate expression with spatial correlation matrix, we first rewrite the channel matrix in a compact form
$\boldsymbol{H}(\boldsymbol{t})=\boldsymbol{\bar{H}}(\boldsymbol{t})[\boldsymbol{\varOmega}(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-1}]^{\frac{1}{2}}+\boldsymbol{S}^{\frac{1}{2}}(\boldsymbol{t})\boldsymbol{U}(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-\frac{1}{2}},$ (71)
where $\boldsymbol{U}=[\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{M}]$ is the Gaussian random variable matrix for NLoS components. Accordingly, $\boldsymbol{H}^{H}\left(\boldsymbol{t}\right)\boldsymbol{H}\left(\boldsymbol{t}\right)$ follows a non-central Wishart distribution [40],
$\boldsymbol{H}^{H}\left(\boldsymbol{t}\right)\boldsymbol{H}\left(\boldsymbol{t}\right)\sim\mathcal{W}_{M}\left(N,\boldsymbol{P},\boldsymbol{\varSigma}\right),$ (72)
with $N$ degrees of freedom, covariance matrix $\boldsymbol{\varSigma}$, and matrix of noncentrality parameters $\boldsymbol{P}$
$\boldsymbol{\varSigma}\triangleq\left(\boldsymbol{\varOmega}+\boldsymbol{I}_{M}\right)^{-\frac{1}{2}}\boldsymbol{S}\left(\boldsymbol{t}\right)\left(\boldsymbol{\varOmega}+\boldsymbol{I}_{M}\right)^{-\frac{1}{2}},$ (73)
$\boldsymbol{P}\triangleq\boldsymbol{\varSigma}^{-1}[\boldsymbol{\varOmega}(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-1}]^{\frac{1}{2}}\boldsymbol{\bar{H}}^{H}(\boldsymbol{t})\boldsymbol{\bar{H}}(\boldsymbol{t})[\boldsymbol{\varOmega}(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-1}]^{\frac{1}{2}}.$ (74)
Now, we can approximate $\boldsymbol{H}^{H}\left(\boldsymbol{t}\right)\boldsymbol{H}\left(\boldsymbol{t}\right)$ by a central Wishart distribution with covariance matrix
$\!\!\boldsymbol{\hat{\varSigma}}=(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-\frac{1}{2}}\boldsymbol{S}(\boldsymbol{t})(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-\frac{1}{2}}$
$\!\!+\!\frac{1}{N}[\boldsymbol{\varOmega}(\boldsymbol{\varOmega}+\boldsymbol{I}_{M})^{-1}]^{\frac{1}{2}}\boldsymbol{\bar{H}}^{H}(\boldsymbol{t})\boldsymbol{\bar{H}}(\boldsymbol{t})[\boldsymbol{\varOmega}\left(\boldsymbol{\varOmega}+\boldsymbol{I}_{M}\right)^{-1}]^{\frac{1}{2}}.\!\!\!$ (75)
and the lower bound ergodic rate for user $m$ is expressed as
$\mathcal{R}_{\text{ZF},m}^{lb}(\boldsymbol{t})=\log_{2}\big(1+\frac{p_{m}}{\sigma_{m}^{2}}\frac{\beta_{m}\left(N-M\right)}{\left[\boldsymbol{\hat{\varSigma}}^{-1}\left(\boldsymbol{t}\right)\right]_{mm}}\big).$ (76)
Using an identity covariance matrix can sometimes oversimplify the channel model, especially if antennas could be positioned arbitrarily close (e.g., $\lambda/10$). However, our design explicitly avoids such scenarios through the imposed minimum antenna spacing constraint to justify this simplification. Specifically, we enforce that the inter-antenna spacing is always greater than $\lambda/2$ as shown in the constraint (6c). Under such configurations, the spatial correlation among different antennas is significantly reduced, and from the physics standpoint, the Rayleigh NLoS components across the array would not experience highly correlated scattering. In the simulation results, we also evaluated the performance deviation between the i.i.d. assumption and the correlated case. The ergodic rates containing the spatial correlation are compared with the simplified ones. Simulations with Bessel correlation functions show that the deviation remains limited and not significant with the minimum antenna spacing $\lambda/2$. In our simulation study, we evaluate the ergodic rate expressions that account for spatial correlation with those obtained under the simplified i.i.d. assumption, demonstrate that with a minimum antenna spacing, the deviation can be minimal and not significant.
VI Simulation Results